OpenStudy (anonymous):

A rectangle has a perimeter of 26 feet. Twice its length (l) is one foot more than three times its width (w). What is the length of the rectangle, in feet?

OpenStudy (anonymous):

A. 5 B. 13 C. 8 D. 16

OpenStudy (anonymous):

Perimeter of rectangle: 2L + 2W = 26 L + W = 13 L = (3W + 1)/2 Substitute: (3W+1)/2 + W = 13 Solve for W: W = 5 Substitute: L + 5 = 13 L = 8